OpenStudy (anonymous):

Combine into a single logarithm. 3LOG (X+Y) + 2LOG(X-Y) - LOG(X^2 + Y^2) =

OpenStudy (anonymous):

start with \[\large n\log(A)=\log(A^n)\] to rewrite as \[\large \log((x+y)^3)+\log((x-y)^2)-\log(x^2+y^2)\]

OpenStudy (anonymous):

then \[\log(A)+\log(B)=\log(AB)\] gives \[\log((x+y)^3(x-y)^2)-\log(x^2+y^2)\]

OpenStudy (anonymous):

and finally \[\log(A)-\log(B)=\log(\frac{A}{B})\]gets \[\log(\frac{(x+y)^3(x-y)^2}{x^2+y^2}\]

OpenStudy (anonymous):

Gracias very much !